# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 1.1. The differential inclusions point of view
In other words, up to a global multiplication by a matrix with determinant 1-1, a compact, connected, elliptic set fulfills (1.9). This, combined with Theorem 1.3, allows us to say that, if $w$ is a Lipschitz solution to (1.8) for an elliptic set $K$, then actually (up to passing to a connected component and up to a global change of sign of determinants) we can assume that $K$ fulfills (1.9), hence there is no real difference between questions (Q) and (Q:DI). This represents a weak converse to Theorem 1.1.
As we see from the above results, the theory of $2\times 2$ elliptic differential inclusions is extremely rich, and this wealth of results and techniques will be essential in our paper. Conversely, our results, especially Theorems A-B-C, represent a step towards a complete understanding of Question (Q:DI).
1.2. General results
Our first main theorem asserts that local $W^{1,\infty}$ bounds are stable with respect to local $L^{\infty}$ convergence for equi-Lipschitz solutions of (1.8).
Theorem A.
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (1.9), $(w_{j})_{j}\subset\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ be equi-Lipschitz solutions to (1.8), $x_{j}\in B_{1}$, and assume $w_{j}$ converges locally uniformly in $B_{1}$ to $w_{\infty}$, $x_{j}\to x_{0}\in B_{1}$. Then we have
$\limsup_{j}\operatorname{diam}([Dw_{j}](x_{j}))\leq 2\operatorname{diam}([Dw_{\infty}](x_{0}))\,$,
where the essential range at a point is defined in (1.12).
It is particularly interesting to apply this result to sequences $w_{j}$ obtained from rescalings of a single solution $w$. To that end, it is convenient to set some notation. Given a Lipschitz map $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ and any point $x_{0}\in B_{1}$, we can consider the rescalings
$w_{r,x_{0}}(h)\doteq\frac{w(x_{0}+rh)-w(x_{0})}{r}.$ (1.13)
As $w$ is Lipschitz, this sequence is precompact in $C^{0}(B)$, for any ball $B\subset\mathbb{R}^{2}$. Therefore, we can introduce the set $\mathcal{B}(w)(x_{0})$ of blow-ups of $w$ at $x_{0}$ to be the collection of maps obtained as locally uniform limits of any subsequence extracted from $w_{r,x_{0}}$. By Rademacher’s theorem $w$ is differentiable at a.e. $x_{0}\in B_{1}$. For any such $x_{0}$, $\mathcal{B}(w)(x_{0})=\{L_{Dw(x_{0})}\}$. In general we observe that, for $K$ compact and fulfilling (1.9),
[…]
due to the strong $W^{1,1}_{\operatorname{loc}}$ convergence of the gradients of rescalings provided by Theorem 1.1.
We may apply Theorem A to a sequence of rescalings as in (1.13) of a given solution $w$, and $x_{j}=x_{0}$ for all $j$. In that case we infer, in particular, that if some blowup $w_{\infty}\in\mathcal{B}(w)(x_{0})$ is linear, say $w_{\infty}(h)=Ah$, then $\operatorname{diam}([Dw_{\infty}](x_{0}))=0$, and just by using the definitions, see (1.12), we obtain that
$\forall\varepsilon>0,\exists\delta>0:\;|Dw(x)-A|\leq\varepsilon,\quad\text{ for a.e. }x\in B_{\delta}(x_{0}).$ (1.15)
We can then say that $Dw$ localizes near the matrix $A$ in sufficiently small neighborhoods of $x_{0}$. Furthermore, (1.15) implies that $w_{\infty}$ is the only element of $\mathcal{B}(w)(x_{0})$. Using Theorem A, we see that to obtain the localization property (1.15) we do not need $w_{\infty}$ to be linear in $\mathbb{R}^{2}$, but we only need it to be differentiable at 0, with differential $A$. Indeed, since $Dw_{\infty}\in K$ a.e. in $\mathbb{R}^{2}$, compare (1.14), then we can employ Theorem A on $w_{\infty}$ itself to deduce that (1.15) holds for $w_{\infty}$, which in turn implies that (1.15) holds for $w$. These considerations allow us to define the regular and the singular set of $w$. To keep the definitions as general as possible, we assume $\Omega\subset\mathbb{R}^{2}$ is open and $w\in\operatorname{Lip}(\Omega,\mathbb{R}^{2})$ solves $Dw\in K$ a.e. in $\Omega$, for $K$ fulfilling (1.9).
Definition 1.5.
We let
[…]
The above discussion shows that, thanks to Theorem A, $x\in\operatorname{Reg}(w)$ if and only if $\mathcal{B}(w)(x)$ consists of a single linear map, that is, $w$ is differentiable at $x$. If $\operatorname{Reg}(w)$ contains an open ball $B$, then we infer that $w$ is $C^{1}$ on $B$, see Lemma 2.1.1, whence the name regular points. We cannot show that $\operatorname{Reg}(w)$ is open in this generality, and we will need to restrict ourselves to cases where $K$ has more structure. Nonetheless, our second result still holds for any elliptic inclusion set $K$, and concerns the size of $\operatorname{Sing}(w)$. We denote by $\mathcal{H}^{\alpha}$ the $\alpha$-dimensional Hausdorff measure.
Theorem B.
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (1.9) and $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve (1.8). Then $\mathcal{H}^{1}(\operatorname{Sing}(w))=0.$
Finally, we will show another remarkable property of solutions to general elliptic differential inclusions. Recall that the Hessian determinant of a smooth function $u$ of two variables can be written as $\det(D^{2}u)=\operatorname{div}\operatorname{div}(Du^{\perp}\otimes Du^{\perp})$, see (5.1), and the latter expression makes sense as a distribution if $u$ is merely $H^{1}$.
Theorem C.
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (1.9) and $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve (1.8). Then, $\det(D^{2}u)$ is a non-positive measure $\mu$. Moreover, if $w_{\infty}=(u_{\infty},v_{\infty})\in\mathcal{B}(w)(x_{0})$, then $\det(D^{2}u_{\infty})=\mu(\{x_{0}\})\delta_{0}$.
Similar properties hold for infinity harmonic functions [33, Theorem 1.5]. Let us add a few words on Theorem C. Ideally, one could hope to deduce from it that $w_{\infty}\in\mathcal{B}(w)(x_{0})$ is linear at points where $\mu(\{x_{0}\})=0$ and that $w_{\infty}$ is one-homogeneous at points where $\mu(\{x_{0}\})\neq 0$. This would tell us that $\operatorname{Sing}(w)$ is actually countable, since a finite measure can have at most countably many atoms. The reason for this hope is that the equation $\det(D^{2}u)=0$ in an open set $\Omega$ implies a very rigid structure of $Du$: essentially, that $Du$ is constant along segments connecting the boundary. This, in turn, yields that $Du$ is constant if $u$ is defined in $\mathbb{R}^{2}$. Such rigidity property has been shown assuming $u\in C^{2}$ [26], $u\in W^{2,\infty}$ [31, §2.6], $u\in W^{2,2}$ [49] and $u$ in the class MA [29]. In all of these cases, $Du$ is (at least weakly) differentiable. Here, however, the only way we can interpret the distribution $\det(D^{2}u)$ is in the very weak sense, see the distribution $\mathcal{D}(u,u)$ in (5.1). The equation $\mathcal{D}(u,u)=0$ is extremely flexible, even if $u\in C^{1,\alpha}$ for $\alpha>0$, see [38, 7, 6]. In a sense, a function $u$ which satisfies $\mathcal{D}(u,u)=0$ and also (1.1) for a strictly monotone $G$, lies in a realm between the known rigidity and flexibility statements. The fact that some rigidity can be expected can be seen in Proposition 7.2, where Theorem C will be instrumental to exclude some types of blowups.
1.3. Regularity for inclusions in elliptic curves
Assume now that $K=\Gamma\subset\mathbb{R}^{2\times 2}$ satisfies (1.9) and is a $C^{1}$ curve, that is, a connected, compact $C^{1}$ submanifold of $\mathbb{R}^{2\times 2}$ of dimension one, which may have a nonempty boundary. Hence it is given by $\Gamma=\gamma(I)$ for some $I=[a,b]$ or $\mathbb{R}/L{\mathbb{Z}}$ and $\gamma\in C^{1}(I,\mathbb{R}^{2\times 2})$ a homeomorphism onto $\Gamma$ with $|\gamma^{\prime}|>0$ on $I$. As previously mentioned, the next Theorem generalizes the main result of [37], where the same conclusion was obtained under the assumption that (1.9) is satisfied with $\sigma(t)=ct^{4}$ for some $c>0$.
Theorem D.
Let $\Omega\subset\mathbb{R}^{2}$ be open. If $w\in\operatorname{Lip}(\Omega,\mathbb{R}^{2})$ satisfies $Dw\in\Gamma$ a.e. in $\Omega$, then $\operatorname{Sing}(w)$ is locally finite. If $\Omega$ is convex, then $\operatorname{Sing}(w)$ contains at most two points. Moreover, $\operatorname{Sing}(w)$ is empty if one of the following sufficient conditions is satisfied.
  • $\Gamma$ is simply connected, that is, $I=[a,b]$;
We also show that the map $Dw$ is locally constant along characteristic lines outside its singular set, see Proposition 6.4, and that the singularities have a rigid structure, see Proposition 6.6.
1.4. Regularity for solutions to degenerate equations
Following [20, 35], for a strictly monotone $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ we define its degeneracy set $\mathcal{D}\doteq\mathcal{D}_{-}\cap\mathcal{D}_{+}$ which can be interpreted as the set of points where the symmetric parts of $DG$ and $(DG)^{-1}$ both have zero eigenvalues. More precisely, the sets $\mathcal{D}_{\pm}$ are:
The latter is also equivalent to $G(X)\in\mathcal{D}_{-}^{c}(G^{-1})$, justifying its interpretation as local ellipticity from above. In [34], it is shown that blow-up limits of solutions to (1.1) are either affine or take values into $\mathcal{D}$. Combining this remarkable fact with the previous Theorems, we obtain a new partial regularity result under structural assumptions on the connected components of $\mathcal{D}$.
Theorem E.
Assume that each connected component of $\mathcal{D}=\mathcal{D}_{-}\cap\mathcal{D}_{+}$ is contained in the first-row projection of a $C^{1}$ curve in $\mathbb{R}^{2\times 2}$. If all but a finite number of these components are
This result is sharp, namely solutions can indeed develop singularities, as shown in [35, Theorem 1.5]. Note that Theorem E applies to the field $G$ constructed in [35, Theorem 1.5], for which $\mathcal{D}$ is the first-row projection of a smooth curve, hence all solutions of (1.1) have a discrete singular set. It is also interesting to apply Theorem E in the variational case $G=Df$ with $f$ strictly convex and radial, that is, $f(x)=g(|x|)$ for some increasing and strictly convex $C^{1}$ function $g\colon[0,\infty)\to[0,\infty)$ such that $g(0)=g^{\prime}(0)=0$. Then we have $X\in\mathcal{D}$ if and only if $|X|\in D_{g}$, where
$D_{g}=\bigcap_{\lambda>0}\overline{\Big\{r\geq 0\colon\liminf_{h\to 0}\frac{g^{\prime}(|r+h|)-g^{\prime}(r)}{h}\leq\lambda\Big\}}\cap\bigcap_{\lambda>0}\overline{\Big\{r\geq 0\colon\limsup_{h\to 0}\frac{g^{\prime}(|r+h|)-g^{\prime}(r)}{h}\geq\frac{1}{\lambda}\Big\}}.$
If we assume that $D_{g}$ is totally disconnected, then all connected components of $\mathcal{D}$ are circles and we deduce that any Lipschitz minimizer of (1.3) is $C^{1}$.
Structure of the paper
In §2 we will introduce the notation, and recall some well-known facts about the essential range that we will need in the sequel. Importantly, we will also explain how to translate elliptic equations into differential inclusions and nonlinear Beltrami systems, and viceversa. Next, §3-4-5-6-7 contain the proofs of Theorems A-B-C-D-E, respectively. Some sections are lengthier than others: in that case we will explain the structure of the section at the start of it.
Acknowledgments
XL is supported by the ANR project ANR-22-CE40-0006.
2. Preliminaries
@@ 2.1. General notation and some elementary facts
$|E|$ denotes the Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$. We write $\mathbf{1}_{E}$ for the indicator function of $E$, namely $\mathbf{1}_{E}(x)=1$ if $x\in E$ and $\mathbf{1}_{E}(x)=0$ otherwise. In the case of super-level sets of a function $u:F\subset\mathbb{R}^{n}\to\mathbb{R}$ (and analogously for sub-level sets and for preimages of intervals) we will use expressions such as $\mathbf{1}_{\{u>\alpha\}}$ to denote the function that, at $x\in F$, returns $1$ if $u(x)>\alpha$ and 0 otherwise. Let $|E|<+\infty$, then $\fint_{E}f(x)dx$ denotes the average of $f$ over $E$. In most cases we will consider only $L^{p}$ spaces with respect to the Lebesgue measure. If we need to say that $\Phi\in L^{p}$ with respect to another finite measure, we will write $L^{p}(\Omega,X;\mu)$, the space of $L^{p}$ functions with respect to the measure $\mu$ on $\Omega$ and with values in the set $X$.
@@
which mean that $a$ is the conformal part of $A$ and $b$ is the anti-conformal part of $A$.
Finally, if $\Omega\subset\mathbb{C}$ is an open set and $f:\Omega\to{\mathbb{C}}$, whenever its differential makes sense we can set
@@ 2.1.1. Essential range and generalized differentials
This, together with (1.12), is closely related to the set of reachable gradients of $w$, see for instance [5, Definition 3.1.10], which is typically defined by considering the set $E$ of differentiability points of $w$. In the case we are interested in, namely for solutions $w$ to (1.8), there is no difference between the two sets: if $K$ fulfills (1.9) and $w$ solves (1.8), then $x_{0}$ is a differentiabilty point for $w$ if and only if $x_{0}$ is a Lebesgue point for $Dw$. Finally, let us show a simple but crucial property of the set $\operatorname{Reg}(w)$, which justifies its name.
Lemma 2.1.
Let $K\subset\mathbb{R}^{2\times 2}$ fulfill (1.9). Let $w$ solve (1.8) and let $\operatorname{Reg}(w)$ be as in Definition 1.2. If $\operatorname{Reg}(w)$ contains a ball $B$, then $w\in C^{1}(B,\mathbb{R}^{2})$.
Proof.
By Theorem A and the discussion preceding Definition 1.2, for all $x_{0}\in\operatorname{Reg}(w)$ we have $Dw](x_{0})=\{A(x_{0})\}$, for some $A(x_{0})\in\mathbb{R}^{2\times 2}$. But then the blow-ups at $x_{0}$, $w_{r,x_{0}}$, actually converge in $W^{1,\infty}_{\operatorname{loc}}$ to $L_{A(x_{0})}$, as can be seen by using (1.12). We infer that $w$ is differentiable everywhere in $B$. Then, continuity follows, again, from (1.12), as $Dw](x_{0})=\{A(x_{0})\}$ implies that for all $\varepsilon>0$, there exists $r>0$ such that $Dw](B_{r}(x_{0}))\subset B_{\varepsilon}(A(x_{0}))=B_{\varepsilon}(Dw(x_{0})).$
@@ 2.2. Three equivalent viewpoints
Let $u\in\operatorname{Lip}(B_{1})$ be a solution to (1.1), where $G$ is defined on $F\supseteq[Du]$. Then, there exists $v\in\operatorname{Lip}(B_{1})$ such that $w\doteq(u,v)$ solves
@@
Moreover, $G$ satisfies (1.5) for all $(x,y)\in F\supseteq[Du]$ with a function $\sigma$ if and only if $K$ fulfills (1.9) with the same $\sigma$. Conversely, if $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ satisfies (1.8) and $K$ fulfills (1.9), then there exists a field $G\in C^{0}(\pi_{1}(K),\mathbb{R}^{2})$ fulfilling (1.5) such that $u$ solves (1.1). Here, $\pi_{1}$ is the projection onto the first row.
@@
if $X_{i},Y_{i}$ are the $i$-th row of $X$ and $Y$ respectively. We thus see that the map $Z\in K\mapsto Z_{1}\in\mathbb{R}^{2}$ is injective. Therefore, it must be invertible when the target is restricted to $\pi_{1}(K)$. Let $I(Z_{1})$ be such inverse, and define also $\tilde{G}(Z_{1})\doteq\pi_{2}(I(Z_{1}))$. Finally, set $G(x,y)\doteq J\tilde{G}(x,y)$, for all $(x,y)\in\pi_{1}(K)$. By definition,
$G(Du)=J\tilde{G}(Du)=JDv,$
so that $u$ solves (1.1). From (2.7), (1.5) also follows, and the proof is concluded. ∎
In the same spirit, as observed in [55, Theorem 3.2], we have:
@@
for $h:F\supseteq[\partial_{z}w]\to\mathbb{C}$. Using notation (2.4), if $K\doteq\left\{(a,h(a)):a\in F\right\}$, then $w$ solves (1.8) for such $K$. Moreover, if $h$ is strictly contractive, in the sense that there exists
@@
then $K$ fulfills (1.9) for some function $\sigma$ with properties (1.6). Conversely, let $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve (1.8), and assume $K$ is a compact set fulfilling (1.9). Then, there exists a strictly contractive $h:[K]_{\mathcal{H}}\to\mathbb{C}$ such that $w$ solves (2.8). We denoted by $K]_{\mathcal{H}}\doteq\{[X]_{\mathcal{H}}:X\in K\}$.
@@
The last viewpoint we introduced in Proposition 2.2, the one of nonlinear Beltrami systems, is another rather useful one, see [22, 2] and references therein. We exploit it now to show how all of our problems (1.1)-(1.8) and (2.8) can be extended in a suitable way so that, for instance, instead of working with $G$ defined on $[Du]$ in (1.1) we can work with $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ fulfilling some special properties at infinity. Thanks to Propositions 2.2-2.2, it is enough to present this extension for nonlinear Beltrami systems.
@@
The existence of a contractive extension $H_{1}$ of $h$ follows from [13, Theorem 3.1]. Then it suffices to set $H=H_{1}\circ\Phi$, where $\Phi\colon{\mathbb{C}}\to{\mathbb{C}}$ is 1-Lipschitz, is the identity on $B_{R}$ for some $R>0$ such that $F\subset B_{R}$, and has compact support. To obtain such a map $\Phi$, one can set for instance $\Phi(z)=\chi(|z|)z$ with
@@
3. Theorem A: elliptic Curves and separation
3.1. Setup and reduction
This section is devoted to the proof of Theorem A. To show it, we need to make a few simplifications. Firstly, we reduce it to the following, cleaner, upper semicontinuity property:

### v2

@@ 2. The differential inclusions point of view
In other words, up to a change of orientation, a compact, connected, elliptic set fulfills (2.4). This, combined with Theorem 2.4, yields that, if $w$ is a Lipschitz solution to (2.3) for an elliptic set $K$, then (up to passing to a connected component) we can assume that $K$ fulfills (2.4). Thus there is no real difference between Questions (Q) and (Q:DI). This represents a weak converse to Theorem 2.2.
3. Notation and preliminaries
@@ 3.1. General notation and some elementary facts
$|E|$ denotes the Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$. We write $\mathbf{1}_{E}$ for the indicator function of $E$, namely $\mathbf{1}_{E}(x)=1$ if $x\in E$ and $\mathbf{1}_{E}(x)=0$ otherwise. In the case of super-level sets of a function $u:F\subset\mathbb{R}^{n}\to\mathbb{R}$ (and analogously for sub-level sets and for preimages of intervals) we will use expressions such as $\mathbf{1}_{\{u>\alpha\}}$ to denote the function that, at $x\in F$, returns $1$ if $u(x)>\alpha$ and 0 otherwise. If $|E|<+\infty$, $\fint_{E}f(x)dx$ denotes the average of $f$ over $E$. In most cases we will consider only $L^{p}$ spaces with respect to the Lebesgue measure. Otherwise, we will write $L^{p}(\Omega,X;\mu)$ for the space of $L^{p}$ functions with respect to the measure $\mu$ on $\Omega$ and with values in the set $X$.
@@
which mean that $a$ is the conformal part of $A$ and $b$ is the anti-conformal part of $A$. Finally, if $\Omega\subset\mathbb{C}$ is an open set and $f:\Omega\to{\mathbb{C}}$, whenever its differential makes sense we can set
@@ 3.1.1. Essential range and generalized differentials
This, together with (2.7), is closely related to the set of reachable gradients of $w$, see for instance [5, Definition 3.1.10], which is typically defined by considering the set $E$ of differentiability points of $w$. In our setting, namely for solutions $w$ to (2.3), there is no difference between the two sets: if $K$ fulfills (2.4) and $w$ solves (2.3), then the set of differentiability points for $w$ and the set of Lebesgue points for $Dw$ coincide.
@@ 3.2. Three equivalent viewpoints
Let $u\in\operatorname{Lip}(B_{1})$ be a solution to (1.1), where $G$ is defined on a compact set $F\supseteq[Du]$. Then, there exists $v\in\operatorname{Lip}(B_{1})$ such that $w\doteq(u,v)$ solves
@@
Moreover, $G$ satisfies (2.1) for all $(x,y)\in F\supseteq[Du]$ for a function $\sigma$ fulfilling (2.2) if and only if $K$ fulfills (2.4) for some $\tilde{\sigma}$ fulfilling (2.2). Conversely, if $w=(u,v)\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ satisfies (2.3) and $K$ is a compact set fulfilling (2.4), then there exists a field $G\in C^{0}(\pi_{1}(K),\mathbb{R}^{2})$ fulfilling (2.1) such that $u$ solves (1.1). Here, $\pi_{1}$ is the projection onto the first row.
@@
if $X_{i},Y_{i}$ are the $i$-th row of $X$ and $Y$ respectively. Through (2.2) we infer that the map $Z\in K\mapsto Z_{1}\in\mathbb{R}^{2}$ is injective. Therefore, it must be invertible when the target is restricted to $\pi_{1}(K)$. Let $I(Z_{1})$ be such inverse, and define also $\tilde{G}(Z_{1})\doteq\pi_{2}(I(Z_{1}))$. Finally, set $G(x,y)\doteq J\tilde{G}(x,y)$, for all $(x,y)\in\pi_{1}(K)$. Writing (3.7) for $X_{1}=a,Y_{1}=b\in\pi_{1}(K)$, we deduce that
$(G(b)-G(a),b-a)\geq\sigma(\sqrt{|b-a|^{2}+|G(b)-G(a)|^{2}}).$
Observing that $G$ is bounded since $K$ is, (2.2) implies that $G$ is continuous on $\pi_{1}(K)$ and that $G$ satisfies (2.1). Moreover, by definition, $G(Du)=J\tilde{G}(Du)=JDv$ a.e., so $u$ solves (1.1), and the proof is concluded. ∎
In the same spirit, as observed in [57, Theorem 3.2], we have:
@@
for $h\colon F\supseteq[\partial_{z}w]\to\mathbb{C}$, and $F$ a compact set. Using notation (3.4), if $K\doteq\left\{(a,h(a)):a\in F\right\}$, then $w$ solves (2.3) for such $K$. Moreover, if $h$ is strictly contractive, in the sense that there exists
@@
then $K$ fulfills (2.4) for some function $\sigma$ with properties (2.2). Conversely, let $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve (2.3), and assume $K$ is a compact set fulfilling (2.4). Then, there exists a strictly contractive $h\colon[K]_{\mathcal{H}}\to\mathbb{C}$ such that $w$ solves (3.8). We denoted by $K]_{\mathcal{H}}\doteq\{[X]_{\mathcal{H}}:X\in K\}$.
@@
The last viewpoint we introduced in Proposition 3.2, the one of nonlinear Beltrami systems, is another rather useful one, see [24, 2] and references therein. We exploit it now to show how all of our problems (1.1)-(2.3) and (3.8) can be extended in a suitable way so that, for instance, instead of working with $G$ defined on $[Du]$ in (1.1) we can work with $G\in C^{0}(\mathbb{R}^{2},\mathbb{R}^{2})$ fulfilling some special properties at infinity. Thanks to Propositions 3.2-3.2, it is enough to present this extension for nonlinear Beltrami systems.
@@
The existence of a contractive extension $H_{1}$ of $h$ follows from [15, Theorem 3.1]. Then it suffices to set $H=H_{1}\circ\Phi$, where $\Phi\colon{\mathbb{C}}\to{\mathbb{C}}$ is 1-Lipschitz, is the identity on $B_{R}$ for some $R>0$ such that $F\subset B_{R}$, and has compact support. To obtain such a map $\Phi$, one can set for instance $\Phi(z)=\chi(|z|)z$ with
@@
4. Elliptic inclusions and separation
This section is devoted to showing the following result, for which it is useful to recall definition (2.7).